Find the perimeter of a triangle if the three sides are:
square root 5 - square root 3
square root 5 + square root 3
4
thanks,

Answer by checkley75(3666)

sqrt5-sqrt3+sqrt5+sqrt3+4=2.236-1.732+2.236+1.732+4=4.472=4=8.472 answer.